[Hallc_running] [New Logentry] RC: Daily update

mdiefent at jlab.org mdiefent at jlab.org
Sun Apr 15 10:50:02 EDT 2018

Logentry Text:
<p><strong>CEBAF status</strong>:</p><ul><li>BLM HV problems during the day shift of April 14 (https://logbooks.jlab.org/entry/3560723). Only 4.0 hours of beam time. </li><li>Various accelerator problems during the swing shift of April 14 (https://logbooks.jlab.org/entry/3560860). Only 4.9 hours of beam time. </li><li>Path length adjustment during the owl shift of April 15 (https://logbooks.jlab.org/entry/3561007). Only 4.5 hours of beam time. </li><li>On Monday, April 16, there will be beam studies for about eight hours. The beam studies will start at 8:00 a.m.. We have been asked to remain locked down for the full duration of the beam studies to allow for tuning for high beam currents. However, we would like to use part of the time window to further investigate the problems with the HMS dipole power supply (in addition to Tuesday). On Saturday evening, we have asked the PD to find out to what extent the other Halls could be included in the beam studies to allow for c
 ontrolled access to Hall C. </li><li>RF recovery is scheduled for Tuesday, April 17, for about six hours. Controlled accesses will be possible during that time period.</li></ul><p><strong>Hall C status</strong>:</p><ul><li>Hall C is configured for the SIDIS experiment (E12-09-017). The HMS is setup for detecting electrons and the SHMS is setup for detecting charged pions and kaons. Details on the experimental setup are available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan#Experimental_setup </li><li>For various momentum and theta settings, the beam current is reduced significantly to keep the pTRIG1 rate below 500kHz. </li><li>Continued problems with the power supply of the HMS dipole during day and swing shifts: The system went into local mode (https://logbooks.jlab.org/entry/3560704, https://logbooks.jlab.org/entry/3560737). Controlled access by Jack Segal with phone support by Steve Lassiter to investigate the problem (https://logbooks.jlab.org
 /entry/3560817). Stable operation of the HMS!
  dipole since then. But we need more time to further investigate and understand the problems with the HMS dipole power supply. </li><li>DAQ problems during the first two hours of the owl shift. Brad Sawatzky replaced the trigger interface card of ROC01 (https://logbooks.jlab.org/entry/3560912). </li></ul><p><strong>Data taking</strong>:</p><ul><li>The detailed run plan is available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan </li><li>For kinematic setting 1, we have completed 10 out of 14 measurements with positive SHMS polarimetry and 2 out of 14 measurements with negative SHMS polarimetry. </li><li>We have lost about four hours ofrun time due to the contineous problems with the HMS dipole power supply and the DAQ (see Hall C status). </li><li>The data taking of the SIDIS experiment with various configuration changes between the runs requires three persons on shift. We have many open slots in the shift schedule (https://misportal.j
 lab.org/mis/physics/shiftSchedule/?experimentRunId=HALLC-2017-2018), in particular for next week’s owl shift. We need more volunteers for shifts and have reached out to all members of the SIDIS/Kaon-LT rungroup. </li></ul><p><strong>Hall C Daily RC meeting</strong>: The will be no meeting today.</p>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3561116
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180415/2d8c97bc/attachment.html>

More information about the Hallc_running mailing list